# How polarization fields can be beneficial in nitride emitters

### v2

@@ Abstract
We revisit the prevailing notion that spontaneous and piezoelectric polarization fields are inherently detrimental to light emission in polar semiconductor heterostructures. Here, we demonstrate that polarization can actually enhance the overlap of electron and hole wave functions in III-nitride quantum wells, which are being considered for polychromatic micron-scale pixels in extended-reality and biomedical applications. By systematically exploring the design space of AlInGaN heterostructures with machine learning, we uncover that larger electric fields globally correlate with higher wave-function overlaps for red emitters. The key is the quantum-confined Stark effect, which enables a reduction in well width while avoiding the associated increase in In composition. Our work demonstrates the use of machine learning to reveal unexpected trends for informing the design of semiconductor heterostructures.
@@
The drop in efficiency of InGaN LEDs with increasing wavelength is attributed to the field-induced separation of electron and hole wave functions in their quantum wells [Fig. 1a].Auf der Maur et al. (1 15); Lin et al. (2023) Achieving red emission requires high In composition, which increases internal electric fields and reduces radiative recombination, quantified by the $B$ coefficient [Fig. 1b]. Unfortunately, making the quantum wells thinner, to bring electrons and holes closer, shifts the emission to shorter wavelengths, necessitating even higher In compositions to maintain red emission [Fig. 1c].
@@
Here, we identify red-emitting quantum wells with the highest wave-function overlaps, by systematically exploring the configuration space of AlxIn1-x-yGayN quantum wells, restricting the search to realistic compositions of 30% or less In. We addressed the combinatorial challenge of such an exploration through ensemble machine-learning (ML) algorithms that provide a 340$\times$ speedup over quantum-mechanical calculations, reproducing a wide range of experimental measurementsDavid et al. (2019); Xue et al. (2023) without any adjustable parameter.
The goal of this search is not to focus on the engineering challenge of optimizing LEDs, which depends sensitively on the consideration of specific structural details. Instead, using idealized models, we take a panoramic view and aim to identify overarching design principles that cannot be discovered through laborious experiments limited to exploring a small region of the design space.
To our surprise, increasing the electric field increases the wave-function overlap in red-emitting quantum wells, as the QCSE enables a reduction in well width while keeping the emission wavelength constant. The wave-function overlaps of the identified red emitters are comparable to state-of-the-art blue emitters. We demonstrate that strategies experimentally shown to improve the performance of red emitters, such as incorporating Al into the barrier and relaxing strain within the quantum well, Hwang et al. (6 23); Ozaki, Funato, and Kawakami (2018); Iida et al. (2020); Li et al. (2023a) also enhance electric fields, raising the possibility that state-of-the-art LED designs may already benefit from increased electric fields.
@@
Using the trained surrogate models, we exhaustively searched the configuration space, screening for optical band gaps between 1.775 eV (698 nm) and 1.925 eV (644 nm), which corresponds to red emission. Initially, we screened for large oscillator strengths by filtering for wave-function squared overlaps that are at least 3$\times$ larger than reference 3 nm wide In0.3Ga0.7N/GaN well. Explicit verification with quantum-mechanical calculations yielded 6591 candidates in total. Out of this pool, 851 candidates feature wave-function squared overlaps at least 5$\times$ larger than the reference; we hereafter refer to these 851 designs as "optimal."
@@
Despite their extreme fields, the wave-function overlaps in the optimal wells surpass those of the reference well with GaN barriers, as illustrated in Fig. 3b. Our analysis also suggests that the optimal quantum wells exhibit $B$ coefficients around $\sim 10^{-11}$ cm3/s, comparable to state-of-the-art blue emitters.Li et al. (2023b) Although the Al-rich barriers confine carriers more strongly within the well because of their larger band offsets,Vichi et al. (2020); Xue et al. (2023) it is the polarization field rather than improved carrier confinement that causes the primary improvement in wave-function overlap. We discuss our numerical control experiments to probe this hypothesis in Section 4 of the SI and Fig. S2.
As the final verification, we push this concept to its logical extreme by considering In0.3Ga0.7N wells with AlN barriers, a combination that maximizes the internal field within our design space. Detailed calculations reveal a field strength of 13 MV/cm and an optical band gap of 1.82 eV at a well width of 1.62 nm. The wave-function overlap associated with this configuration is indeed an order of magnitude larger than that of 3 nm InGaN/GaN quantum wells, as shown in Fig. 3b. Our initial search space did not encompass well widths of 1.62 nm, reflecting the practical necessity of discretizing continuous variables to numerically represent them. Nevertheless, this limitation did not hinder us from identifying the broader design rule that ultimately led to the discovery of this configuration.
In addition to compositional engineering, there is more freedom in modifying polarization fields by exploring qualitatively different heterostructure designs. We facilitated this exploration by considering InGaN quantum wells separated from GaN barriers with AlInGaN interlayers, a class of heterostructures already employed in state-of-the-art LEDs.Hwang et al. (6 23); Iida et al. (2020); Li et al. (2023a) We recover the limit of quantum wells with pure AlInGaN barriers, corresponding to our initial search space, by setting the interlayer thickness equal to the barrier thickness. To streamline our understanding, we focused on AlN interlayers, although our conclusions apply to other compositions.
@@
Notably, our findings align with experimental reports of the most efficient red LEDs, many of which feature Al-containing barriers or strain-relaxed quantum wells.Hwang et al. (6 23); Iida et al. (2020); Li et al. (2023a) Hence, we suggest critically reevaluating whether polarization fields already play a role in improving the performance of these LEDs. Overall, carriers recombine more quickly in quantum wells with strong light-matter coupling, meaning fewer carriers are needed to achieve high current densities. This helps to reduce the carrier-density-dependent hue shiftPant et al. (2022) and mitigate efficiency droop, a phenomenon where the efficiency decreases with increasing carrier density.Li et al. (2023b)
The extent to which polarization can be further useful in practically designing efficient light emitters remains an open question. We have not explicitly modeled complex effects such as well-width fluctuations, inter-well carrier transport, and device-level characteristics, the predictive modeling of which remains a formidable and ongoing challenge. Moreover, there are significant engineering challenges that large electric fields present in LEDs; we discuss these and possible solutions in Section 7 of the SI.
Nevertheless, we have provided clear evidence that internal fields do not need to be mitigated to enhance wave-function overlap. This is compelling given that polarization fields are the reason why thin InGaN quantum wells ($\leq 3$ nm) with low In concentrations ($\leq$ 30%), used in red LEDs, can emit red light in the first place.
@@
III-nitrides are a cornerstone of optoelectronics. With their distinct crystal structure in the III-V family and their proliferation in industry, they offer an unmatched platform to harness polarization effects. While we do not claim to solve the non-trivial challenges associated with further amplifying electric fields, we believe that the challenges are equally great, if not greater, for designs that mitigate these fields. This is a salient point to consider as the community considers moving towards nonpolar designs.Lin et al. (2023) The ability to exploit the unique polarization degree of freedom in the design of optoelectronics should not be underestimated.
@@ Acknowledgements
We thank Kyle Bushick, Mahlet Molla, Emily Oliphant, and Amanda Wang for their feedback. This project was funded by the U.S. Department of Energy, Office of Energy Efficiency and Renewable Energy, under Award No. DE-EE0009163. Computational resources were provided by the National Energy Research Scientific Computing Center, a Department of Energy Office of Science User Facility, supported under Contract No. DEAC0205CH11231. N.P. acknowledges the support of the Natural Sciences & Engineering Research Council of Canada Postgraduate Scholarship.

### v5

# Enhancing light emission with electric fields in polar nitride semiconductors

@@ Abstract
Significant effort has been devoted to mitigating polarization fields in nitride LEDs, as these fields are traditionally viewed as detrimental to light emission, particularly for red emission. Contrary to this prevailing notion, we demonstrate that strong polarization fields can enhance the optical-transition strength of AlInGaN quantum wells emitting in the red, which has been historically challenging to achieve. By leveraging machine-learning surrogate models trained on multi-scale quantum-mechanical simulations, we globally explore the heterostructure design space and uncover that larger fields correlate with higher electron-hole overlap. This relation arises from the quantum-confined Stark effect, which enables thinner wells without requiring higher indium compositions, thus overcoming a key limitation in nitride epitaxy. Structural and compositional engineering of internal fields offers a unique dimension for designing polychromatic nitride LEDs, crucial for miniaturizing LED pixels to the micron scale for extended-reality and biomedical applications. Broadly, our work demonstrates how machine learning can uncover unexpected paradigms for semiconductor design.
micro-LED, light-emitting diode, quantum well, III-nitride, polarization fields, machine learning, multi-scale modeling, red emission
@@ I Introduction
The drop in efficiency of InGaN LEDs with increasing wavelength is attributed to the field-induced separation of electron and hole wave functions in their quantum wells [Fig. 1a].Auf der Maur et al. (1 15); Lin et al. (2023) This reduces the optical-transition dipole moment, which is proportional to the wave-function overlap. Achieving red emission requires high In composition, which increases internal electric fields and reduces radiative recombination, quantified by the $B$ coefficient [Fig. 1b]. Unfortunately, making the quantum wells thinner, to bring electrons and holes closer, shifts the emission to shorter wavelengths, necessitating even higher In compositions to maintain red emission [Fig. 1c].
@@
Here, we identify red-emitting quantum wells with the highest electron-hole wave-function overlaps, by systematically exploring the configuration space of AlxIn1-x-yGayN quantum wells, restricting the search to realistic compositions of 30% or less In. By taking a panoramic view, we aim to identify overarching design principles that are difficult to access with laborious experiments limited to exploring a small region of the design space. We addressed the combinatorial challenge of such an exploration through ensemble machine-learning (ML) algorithms that provide a 340$\times$ speedup over quantum-mechanical calculations, reproducing a wide range of experimental measurementsDavid et al. (2019); Xue et al. (2023) without any adjustable parameter.
To our surprise, increasing the electric field increases the wave-function overlap in red-emitting quantum wells, as the quantum-confined Stark effect (QCSE) enables a reduction in well width while keeping the emission wavelength constant. The wave-function overlaps of the identified red emitters are comparable to state-of-the-art blue emitters. We propose that an enhancement of electric fields could partially explain why incorporating Al into the barrier or relaxing strain within the quantum well improves the performance of state-of-the-art red LEDs. These strategies have been shown to improve LED performance in several experimental works, Hwang et al. (6 23); Ozaki, Funato, and Kawakami (2018); Iida et al. (2020); Li et al. (2023a) though the beneficial role of polarization fields was not previously identified.
II Machine-learning accelerated search
@@
Using the trained surrogate models, we exhaustively searched the configuration space, screening for optical band gaps between 1.775 eV (698 nm) and 1.925 eV black(644 nm), which corresponds to red emission. Initially, we screened for large oscillator strengths by filtering for wave-function squared overlaps that are at least 3$\times$ larger than reference 3 nm wide In0.3Ga0.7N/GaN well. Manual verification with explicit quantum-mechanical calculations yielded 6591 candidates in total. Out of this pool, 851 candidates feature wave-function squared overlaps at least 5$\times$ larger than the reference; we hereafter refer to these 851 designs as "optimal."
III Optimal Designs
@@
Despite their extreme fields, the wave-function overlaps in the optimal wells surpass those of the reference well with GaN barriers, as illustrated in Fig. 3b. Our analysis also suggests that the optimal quantum wells exhibit $B$ coefficients around $\sim 10^{-11}$ cm3/s, comparable to state-of-the-art blue emitters.Li et al. (2023b) Although the Al-rich barriers confine carriers more strongly within the well because of their larger band offsets,Vichi et al. (2020); Xue et al. (2023) it is the polarization field rather than improved carrier confinement that causes the primary improvement in wave-function overlap (Fig. 3c). We discuss our numerical control experiments to probe this hypothesis in Section 4 of the SI and Fig. S2.
As the final verification, we push this concept to its logical extreme by considering In0.3Ga0.7N wells with AlN barriers, a combination that maximizes the internal field within our design space. Detailed calculations reveal a field strength of 13 MV/cm and an optical band gap of 1.82 eV at a well width of 1.62 nm. The wave-function overlap associated with this configuration is indeed an order of magnitude larger than that of 3 nm InGaN/GaN quantum wells, as shown in Fig. 3b. Our initial search space did not encompass well widths of 1.62 nm, reflecting the practical necessity of discretizing continuous variables to numerically represent them. Nevertheless, this limitation did not hinder us from identifying the broader design rule that ultimately led to the discovery of this configuration. While it is not presently feasible to use pure AlN barriers in actual LEDs, this example illustrates the fact that polarization fields need not be mitigated to enhance the wave-function overlap.
In addition to compositional engineering, there is more freedom in modifying polarization fields by exploring qualitatively different heterostructure designs (explored in Section 5 of the SI). We facilitated this exploration by considering InGaN quantum wells separated from GaN barriers with AlInGaN interlayers, a class of heterostructures already employed in state-of-the-art LEDs.Hwang et al. (6 23); Iida et al. (2020); Li et al. (2023a) We recover the limit of quantum wells with pure AlInGaN barriers, corresponding to our initial search space, by setting the interlayer thickness equal to the barrier thickness. To streamline our understanding, we focused on AlN interlayers, although our conclusions apply to other compositions.
@@ IV Discussion
Notably, our findings align with experimental reports of the most efficient red LEDs, many of which feature Al-containing barriers or strain-relaxed quantum wells.Hwang et al. (6 23); Iida et al. (2020); Li et al. (2023a) Hence, we suggest critically reevaluating whether polarization fields already play a role in improving the performance of these LEDs. Overall, carriers recombine more quickly in quantum wells with higher wave-function overlap, meaning fewer carriers are needed to achieve high current densities. This helps to reduce the carrier-density-dependent hue shiftPant et al. (2022) and mitigate efficiency droop, a phenomenon where the efficiency decreases with increasing carrier density.Li et al. (2023b) We underscore that the most extreme fields are not needed to reap the benefits. Even slight increases in the fields allow quantum wells to be thinner without increasing the In content, thanks to the QCSE.
However, it remains an open question if further enhancing polarization fields will lead to substantial benefits when considering other aspects of LED performance. Notably, we have not explicitly modeled well-width fluctuations, inter-well carrier transport, and device-level characteristics, all of which present ongoing challenges for predictive modeling. Furthermore, Al-containing barriers and large electric fields in LEDs introduce several engineering hurdles. These include the competition between radiative and Auger-Meitner recombination, spectral linewidth broadening, material challenges during growth, and challenges with electrical injection. We discuss these issues and potential solutions in the SI Section 7, including identifying nearly lattice-matched designs provided in Supporting Dataset 2. We also refer the reader to Ref. 40 in which we discuss the challenges of thin quantum wells in the context of green LEDs.
V Conclusion
@@
III-nitrides are a cornerstone of optoelectronics. With their distinct crystal structure in the III-V family and their proliferation in industry, they offer an unmatched platform to harness polarization effects. While we do not claim to solve the challenges associated with further amplifying electric fields, we note that designs mitigating these fields face comparable, if not greater, challenges. This is a salient point to consider as the community considers moving towards nonpolar designs.Lin et al. (2023) The ability to exploit the unique polarization degree of freedom in the design of optoelectronics should not be underestimated, especially with the addition of new elements to the III-nitride familyJena et al. (2019) that introduce ferroelectricity.
We have shown that internal fields do not need to be mitigated to enhance wave-function overlap. This is compelling given that polarization fields are the reason why thin InGaN quantum wells ($\leq 3$ nm) with low In concentrations ($\leq$ 30%), used in red LEDs, can emit red light in the first place. Despite two decades of efforts to mitigate polarization fields, the most efficient visible LEDs, especially in Industry, exhibit fields on the order of MV/cm. Hence, polarization is a double-edged sword, which, despite its drawbacks, also benefits LEDs through the QCSE by enabling long wavelength emission in thin wells with low In concentrations.
Associated Content
Supporting Information. Additional details of the computational methods and parameters used (PDF). Survey of the optimal quantum wells and geometric blueprint for extreme fields (PDF). Discussion of impact of alloy fluctuations, carrier confinement, emission broadening due to well-width and alloy fluctuations, challenges associated with large electric fields, and electron overflow (PDF). Dataset of ideal quantum-well configurations (Dataset 1) and nearly lattice-matched configurations (Dataset 2).
@@ Acknowledgements
N.P. thanks Kyle Bushick, Mahlet Molla, Emily Oliphant, and Amanda Wang for their feedback on the manuscript.
Funding Sources. This project was funded by the U.S. Department of Energy, Office of Energy Efficiency and Renewable Energy, under Award No. DE-EE0009163. Computational resources were provided by the National Energy Research Scientific Computing Center, a Department of Energy Office of Science User Facility, supported under Contract No. DEAC0205CH11231. N.P. acknowledges the support of the Natural Sciences & Engineering Research Council of Canada Postgraduate Scholarship.